OpenStudy (anonymous):

a = 5, b = 12, c = √(119) √(154) 13

OpenStudy (anonymous):

@RockerSk8er

OpenStudy (anonymous):

@lianacarolina do u happen to know?

OpenStudy (anonymous):

I'm working on it. Haha

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] \[5^2 + 12^2 = c^2\] \[25 + 144= c^2\] \[c^2 = 169\] \[\sqrt{169} = 13\]

OpenStudy (anonymous):

thnx:)

OpenStudy (anonymous):

No problem. :)